if A+B=pi/2, where A,B>0, then the maimum value of sinA + sinB is equal to

Grade:11

1 Answers

Suraj Singh
7 years ago
Hy darshini A+B=pi/2... Therefore A=pi/2 - B.. We have to max value of SinA+SinB.. Putting tha above value it comes SinA+CosB whose maximum value is 1/root 2.. Hope U like it..